A quality control engineer takes daily samples of 10 electronic components and checks them for inspections. On 200 consecutive working days, he obtained 112 days with zero defective, 76 days with one defective and 12 days with two defective. If these samples can be looked upon as samples from binomial population, obtain the expected or theoretical number of days with zero defective one defective and 2 or more defective.


1
Expert's answer
2021-03-07T16:54:22-0500

Probability of days with 0 defective component "=\dfrac{112}{200} =0.56"


Probability of days with 1 defective component "=\dfrac{76}{200} =0.38"


Probability of days with 2 defective component "=\dfrac{12}{200} =0.06"


The probability Distribution table is-





So Expected value is

"E(X)=\sum X.P(X)"


"=0+0.38+0.12=0.50"


Hence The expected value is 0.50.